# Quantum-enhanced Markov Chain Monte Carlo for systems larger than your Quantum Computer

### v1

@@ 2.3 Thermalisation
Upon initialisation, the Markov-chain proceeds until thermalisation is achieved, signifying convergence to the stationary distribution. Subsequently, one can then sample from the chain at intervals spaced to ensure uncorrelated samples which is called the auto-correlation time of the Markov Chain. The expectation value of observables, such as magnetisation, can then be estimated from these non-correlated samples. However, the crux lies in thermalisation. A local proposal method, altering one spin at a time, tends to get trapped in local minima. Conversely, a uniform random proposal explores the energy landscape inefficiently, being entirely non-local. In the low temperature regime, both approaches exhibit slow auto-correlation and thermalisation times. It should be noted that the uniform distribution will faithfully converge in a number of steps similar to that of the size of the entire state space. In other words, given the time it would take to solve the problem by brute force, the uniform proposal can approximate the solution. Thus, the uniform proposal provides an upper bound to the possible thermalisation times of MCMC.
@@
One of the issues with thermalisation regards how one can tell when a Markov chain has reached the stationary distribution, and how many steps to take as auto-correlation time. Multiple techniques have been proposed, however many are deemed to be fickle or flawed [12, 13, 14]. For example, it is possible for two different chains on the same problem instance to both individually appear like they have converged - their observables appear stable. When they are compared side by side however, it is clear they they have “converged” to different local distributions as the observables calculated from sampling the chains are completely different [15]. Alternatively, two chains may agree on observables while neither seeming to be in a stationary distribution. The former is a particular problem when updates are particularly local.
@@
Classical techniques such as parallel tempering or population annealing sample multiple different chains, allowing one to parallelise computation and mitigate for the issues of convergence with local proposals [16, 17, 18, 19]. These techniques are all directly applicable to CGQeMCMC, meaning any gain in convergence is propagated easily to state of the art techniques.
2.4 Quantum enhancement
@@
Recently, the QeMCMC has extended the classical MCMC to a Noisy Intermediate Scale Quantum (NISQ) computing setting, with a cubic reduction in MCMC thermalisation time for some problems [1]. In QeMCMC, the proposal states of the Markov chain are the outcome of a quantum computation, with speedup derived from reducing the number of required update steps. The first step is to encode the current state of the Markov Chain as binary input to the computational basis state of the quantum computer - mapping the $\{1,-1\}^{n}$ spin states to $\{1,0\}^{n}$ computational basis states. One then evolves this state according to a Hamiltonian ($H$) .
@@
This unitary evolution takes the form: $U=e^{-iHt}$ with practical implementation requiring an approximation through the Trotter-Suzuki product. The problem Hamiltonian ($H_{prob}$) encodes the classical model instance and although the mixing Hamiltonian could take many different forms, it is defined here as [^5] The notation is kept in line with previous literature on the subject where possible [1, 7] $H_{\mathrm{mix}}=\sum_{j=1}^{n}X_{j}$. Here, hyper-parameters are sampled in the range $t=(2,20)$ and $\gamma=(0.25,0.6)$, although optimisation of these parameters is an ongoing issue.
$H_{\text{prob }}=-\sum_{j>k=1}^{n}J_{jk}Z_{j}Z_{k}-\sum_{j=1}^{n}h_{j}Z_{j}=\sum_{\boldsymbol{s}}E(\boldsymbol{s})|\boldsymbol{s}\rangle\langle\boldsymbol{s}|$ (7)
Measurement in the computational basis state returns a binary outcome that is used as the proposal state as seen in Algorithm 2. The Hamiltonian is inspired by the adiabatic program, and perturbative analysis expects it to propose states similar in energy to the previous state, and far in Hamming distance[1]. Both of these properties are optimal for fast MCMC convergence rate, and cannot be replicated using classical computers. Other choices of unitary evolution have been proposed to replace the time-independent Hamiltonian evolution, however these have not yet been explored experimentally and thus wont be considered here [20].
@@
// Initialisation
2 $\boldsymbol{s}\leftarrow$ initial spin configuration;
@@ 5 $t\leftarrow$ hyper-parameter in range $(2,20)$;
6 $|\psi\rangle=\exp[-iHt]|\boldsymbol{s}\rangle$ ;
7 $\boldsymbol{s}^{\prime}=$ X measurement result of $|\psi\rangle$ ;
@@
The quantum proposal distribution, $Q\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)=\left|\left\langle\boldsymbol{s}^{\prime}|U|\boldsymbol{s}\right\rangle\right|$, has the symmetry requirement: $\left|\left\langle\boldsymbol{s}|U|\boldsymbol{s}^{\prime}\right\rangle\right|=\left|\left\langle\boldsymbol{s}^{\prime}|U|\boldsymbol{s}\right\rangle\right|^{2}$. Normally, one is required to calculate $Q\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)$ explicitly as in Equation 4, however as it is merely the ratio of $\frac{Q\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)}{{Q\left(\boldsymbol{s}\mid\boldsymbol{s}^{\prime}\right)}}$ required, the equation easily simplifies thanks to the symmetry constraint.
$Q\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)=\left|\left\langle\boldsymbol{s}^{\prime}|U|\boldsymbol{s}\right\rangle\right|^{2}=\left|\left\langle\boldsymbol{s}|U|\boldsymbol{s}^{\prime}\right\rangle\right|^{2}=Q\left(\boldsymbol{s}\mid\boldsymbol{s}^{\prime}\right)$ (8)
@@
$A\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)=\min\left(1,\frac{\mu\left(\boldsymbol{s}^{\prime}\right)}{\mu(\boldsymbol{s})}\right)=\min\left(1,\exp{\frac{E(\boldsymbol{s})-E(\boldsymbol{s}^{\prime})}{T}}\right)$ (9)
Of course, this elegant simplification of the Metropolis acceptance criteria is not enough to suggest quantum advantage. Both perturbation theory and numerical analysis have shown that a quantum update as above have some remarkable properties [1]. Proposals that maximise the Hamming distance between $\boldsymbol{s}$ and $\boldsymbol{s}^{\prime}$ (as in the uniform update) while minimising the Energy difference (as in the local update) and are thus likely to be accepted are optimal. This is exactly the criteria that the QeMCMC meets, meaning it has freedom of movement on the lattice at no cost to acceptance probability. One can view this advantage as “searching” over a superposition of possible states of similar energy, before collapsing to one at measurement. This becomes incredibly important when in the low temperature regime. As mentioned in Section 2.2, Markov-chain suffer from low acceptance rate in the low temperature regime so proposals that can suggest similar energy states are highly likely to be accepted and can significantly improve a Markov-chains convergence.
3 Framework and figures of merit
In this section, we establish the foundational framework for employing CGQeMCMC techniques on the Ising model by outlining the key components and methodologies. Additionally, we provide a comprehensive overview of two pivotal figures of merit: the absolute spectral gap and magnetisation estimation. These metrics are introduced to provide background for the numerical experiments detailed in Section 5.
3.1 Framework
In CG modelling, the behaviour of a complex system is simulated by using a simpler representation of the system. This is immediately realised in Ising models at each MCMC step by isolating a group consisting of $q$ spins - here referred to by the corresponding fraction of the lattice per group $(q/n)$. These spins may not necessarily be close to one another or even interact, however the optimal (in terms of thermalisation) choice of spins in a group is likely to depend on the details of the problem. The coarse graining is shown in Algorithm 3, where the QeMCMC (Algorithm 2) is run as a subroutine on a small group $s_{g}$.
@@ 8 $s_{g}^{\prime}\leftarrow$ QeMCMC ($s_{g}$);
// update $\boldsymbol{s}^{\prime}$ with results from $s^{\prime}_{g}$ QeMCMC
9 $\boldsymbol{s}^{\prime}[index]\leftarrow s_{g}^{\prime}$ ;
@@
Of course, there is scope for a multitude of different groupings, sub-groupings and clusterings as is sometimes seen in classical literature [21, 22, 23, 24, 25, 26]. These are not exhaustively discussed here, where we simply give the framework of coarse graining the Ising model lattice along with some illustrative examples that demonstrate the potential of the approach. Discussion of the choice of spins to include in the active group, as well as the parameters chosen as input to the QeMCMC, including the choice of the “reduced” Hamiltonian, are left to Section 4. It should be made clear that any Quantum enhancement found can most likely be combined with traditional MCMC methods in hybrid schemes that are likely to outpace “plain” MCMC, regardless of proposal.
3.2 Spectral Gap
As mentioned in Section 2.3, estimating the thermalisation time $(\tau_{\varepsilon})$ of an algorithm is not simple, as one cannot necessarily tell when the Markov chain is in the stationary state. There is also a dependence on the starting state of the algorithm meaning between successive MCMC runs thermalisation can differ in steps by orders of magnitude. The thermalisation time is however bounded analytically:
@@ 3.2 Spectral Gap
where $\varepsilon$ is the error in total variational distance used to quantify the bounds of $(\tau_{\varepsilon})$ [27]. The Spectral gap $(\delta)$ describes the slowest facet of a chains convergence and is found through the transition probability matrix, $P$. Given the eigenvalues $\{\lambda\}$ of $P$, $\delta$ is calculated:
@@
The Spectral gap is inversely proportional to the upper bound of the thermalisation time (Equation 10) and thus provides an unambiguous quantification of a Markov chains convergence. Numerical calculation of $\delta$ is very inefficient however, requiring numerical integration (often by random sampling, see A) to create the $(2^{n}\times 2^{n})$ $P$ matrix before Eigen-decomposition. Due to this, calculation of the spectral gap is exponentially more demanding than an actual MCMC algorithm.
@@
The spectral gap also ignores the use of a particular chain in more sophisticated schemes mentioned in Section 2.3 that rely on running multiple chains concurrently. Thus, one must be reluctant to overstate the results of spectral gap analysis.
3.3 Convergence analysis
@@
To analyse the convergence of CGQeMCMC in test cases that is currently technologically inaccessible to QeMCMC, we run Markov Chains and estimate the average magnetisation and energy of the Boltzmann distribution. For low temperatures, this essentially becomes an optimisation problem as only a few states contribute.
@@
As mentioned in Section 2, calculating the average magnetisation $\langle m\rangle_{\mu}$ is dependent on first calculating the partition function. This makes direct calculation infeasible for even relatively simple systems. The alternative is to perform the following estimation where $m(s)$ is calculated for each sampled state by Equation 3:
$\langle m\rangle_{\mu}\approx\frac{1}{N}\sum_{\boldsymbol{s}\in\mathcal{S}}m(\boldsymbol{s})$ (12)
@@
Assuming thermalisation of the Markov chain, this is accurate with high probability even when $N$ is very small in comparison to the entire state space [28, 27]. One can use estimation of this observable as a figure of merit by asking the question: “How well can our Markov chain estimate the true magnetisation?”. Of course this requires knowledge of the stationary distribution. Here the stationary distribution of exemplar 16, 25 and 36 spin instances are found by brute force, meaning that the update proposals can be evaluated against the true magnetisation as benchmark.
@@
Specifically, this paper will look at cumulative magnetisation as sampling from Markov Chains generally requires thinning (to take non-correlated samples) and burning (to ensure thermalisation) which incurs hyper-parameters complicating the analysis.
@@
Crucially, in certain cases, magnetisation estimation can be a flawed figure of merit. As briefly discussed in Section 2.3, a chain may appear to converge to the stationary distribution, while it is actually stuck in a local minimum. If the magnetisation contribution of this local minimum is similar to that of the actual magnetisation, then this becomes a flawed figure of merit. Thus, alongside the magnetisation calculation we analyse the convergence of average energy between multiple chains which can be compared with the exact average energy for the stationary distribution. If local minima impede convergence, this measure will display an obvious bias.
4 Coarse graining strategies
@@
In this section, we give three different coarse graining strategies: (a) a “naive” approach where one chooses a random subset of spins and ignores the rest, (b) an improved local approach where one chooses a random subset of spins while treating the remaining spins as an external a field and (c) an approach that one considers multiple subsets, treating each of them separately as in the earlier approaches.
@@
Before proceeding to the proposed strategies, let us highlight two properties of the QeMCMC method that were key to the advantage demonstrated in [1]. The first is that the Hamiltonian used (evolution to obtain the proposal step) includes the problem’s Hamiltonian, and with suitable choice of parameters the proposal step does not differ significantly in energy – thus ensuring proposed steps that are accepted even when being close to thermalisation. The second is that, unlike classical local strategies, QeMCMC can lead to proposal step that differs (significantly) in Hamming distance from the previous step, “spanning/searching” the full space and escaping from local minima.
4.1 Naive CG - local groups
The naive choice of coarse graining is to pick spins that are neighbours forming a local group, $g_{l}={s_{r},s_{r+1}}...s_{r+q}$ where $s_{r}$ is a randomly selected spin that begins the group [^6] We need to stress here that the spins and their “locality” refers to their labels, that in some cases (e.g. 1-d Ising model) also carry physical information, while in other cases (random Ising model) do not. We still use the term “local” to distinguish this coarse graining strategy form others, but we want to make sure that this is not mistakenly understood as “physically” local.. Once isolated, the group is updated according to the QeMCMC routine outlined in Section 2.4 with $\tilde{J}$ and $\tilde{h}$ the parameters defining the “partial Hamiltonian” of the newly formed group. This new Hamiltonian essentially ignores the spins not selected:

### v2

@@ II.3 Thermalization
Upon initialization, the Markov-chain proceeds until thermalization is achieved, signifying convergence to the stationary distribution. Subsequently, one can then sample from the chain at intervals spaced to ensure uncorrelated samples which is called the auto-correlation time of the Markov Chain. The expectation value of observables, such as magnetization, can then be estimated from these non-correlated samples. However, the crux lies in thermalization. A local proposal method, altering one spin at a time, tends to get trapped in local minima. Conversely, a uniform random proposal explores the energy landscape inefficiently, being entirely non-local. In the low temperature regime, both approaches exhibit slow auto-correlation and thermalization times. It should be noted that the uniform distribution will faithfully converge in a number of steps similar to that of the size of the entire state space. In other words, given the time it would take to solve the problem by brute force, the uniform proposal can approximate the solution. Thus, the uniform proposal provides an upper bound to the possible thermalization times of MCMC.
@@
One of the issues with thermalization regards how one can tell when a Markov chain has reached the stationary distribution, and how many steps to take as auto-correlation time. Multiple techniques have been proposed, however many are deemed to be fickle or flawed Vehtari et al. (2021); Flegal (2008); Geyer (2011). For example, it is possible for two different chains on the same problem instance to both individually appear like they have converged - their observables appear stable. When they are compared side by side however, it is clear they they have “converged” to different local distributions as the observables calculated from sampling the chains are completely different Moins et al. (2023). Alternatively, two chains may agree on observables while neither seeming to be in a stationary distribution. The former is a particular problem when updates are local.
@@
Classical techniques such as parallel tempering or population annealing sample multiple different chains, allowing one to parallelize computation and mitigate for the issues of convergence with local proposals Machta and Ellis (2011); Wang et al. (2015); Weigel et al. (2021); Swendsen and Wang (1986). These techniques are all directly applicable to CGQeMCMC, meaning any gain in convergence is propagated easily to state of the art techniques.
II.4 Quantum enhancement
@@
Recently, the QeMCMC has extended the classical MCMC to a Noisy Intermediate Scale Quantum (NISQ) computing setting, with a cubic reduction in MCMC thermalization time for some problems Layden et al. (2023). In QeMCMC, the proposal states of the Markov chain are the outcome of a quantum computation, with speedup derived from reducing the number of required update steps. The first step is to encode the current state of the Markov Chain as binary input to the computational basis state of the quantum computer - mapping the $\{1,-1\}^{n}$ spin states to $\{1,0\}^{n}$ computational basis states. One then evolves this state according to a Hamiltonian ($H$) .
@@
This unitary evolution takes the form: $U=e^{-iHt}$ with practical implementation requiring an approximation through the Trotter-Suzuki product. The problem Hamiltonian ($H_{prob}$) encodes the classical model instance and although the mixing Hamiltonian could take many different forms, it is defined here as $H_{\mathrm{mix}}=\sum_{j=1}^{n}X_{j}$  [^3] The notation is kept in line with previous literature on the subject where possible Layden et al. (2023); Nakano et al. (2023). Here, hyper-parameters are sampled in the range $t=(2,20)$ and $\gamma=(0.25,0.6)$, although optimization of these parameters is an ongoing issue.
$H_{\text{prob }}=-\sum_{j>k=1}^{n}J_{jk}Z_{j}Z_{k}-\sum_{j=1}^{n}h_{j}Z_{j}=\sum_{\mathbf{s}}E(\mathbf{s})|\mathbf{s}\rangle\langle\mathbf{s}|$ (7)
Measurement in the computational basis state returns a binary outcome that is used as the proposal state as seen in Algorithm 2. The Hamiltonian is inspired by the adiabatic program, and perturbative analysis expects it to propose states similar in energy to the previous state, and far in Hamming distance Layden et al. (2023). Both of these properties are optimal for fast MCMC convergence rate, and cannot be replicated using classical computers. Other choices of unitary evolution have been proposed to replace the time-independent Hamiltonian evolution, however these have not yet been explored experimentally and thus wont be considered here Mazzola (2021).
@@
// Initialization
2 $\mathbf{s}\leftarrow$ initial spin configuration;
@@ 5 $t\leftarrow$ hyper-parameter in range $(2,20)$;
6 $|\psi\rangle\leftarrow\exp[-iHt]|\mathbf{s}\rangle$ ;
7 $\mathbf{s}^{\prime}\leftarrow$ X measurement result of $|\psi\rangle$ ;
@@
The quantum proposal distribution, $Q\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)=\left|\left\langle\mathbf{s}^{\prime}|U|\mathbf{s}\right\rangle\right|$, has the symmetry requirement: $\left|\left\langle\mathbf{s}|U|\mathbf{s}^{\prime}\right\rangle\right|=\left|\left\langle\mathbf{s}^{\prime}|U|\mathbf{s}\right\rangle\right|$. Normally, one is required to calculate $Q\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)$ explicitly as in Equation 4, however as it is merely the ratio of $\frac{Q\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)}{{Q\left(\mathbf{s}\mid\mathbf{s}^{\prime}\right)}}$ required, the equation easily simplifies thanks to the symmetry constraint.
$Q\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)=\left|\left\langle\mathbf{s}^{\prime}|U|\mathbf{s}\right\rangle\right|^{2}=\left|\left\langle\mathbf{s}|U|\mathbf{s}^{\prime}\right\rangle\right|^{2}=Q\left(\mathbf{s}\mid\mathbf{s}^{\prime}\right)$ (8)
@@
$A\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)=\min\left(1,\frac{\mu\left(\mathbf{s}^{\prime}\right)}{\mu(\mathbf{s})}\right)=\min\left(1,\exp{\frac{E(\mathbf{s})-E(\mathbf{s}^{\prime})}{T}}\right)$ (9)
Of course, this elegant simplification of the Metropolis acceptance criteria is not enough to create a useful quantum algorithm. Both perturbation theory and numerical analysis have shown that a quantum update as above have some remarkable properties Layden et al. (2023). Proposals that maximize the Hamming distance between $\mathbf{s}$ and $\mathbf{s}^{\prime}$ (as in the uniform update) while minimising the Energy difference (as in the local update) and are thus likely to be accepted are optimal. This is exactly the criteria that the QeMCMC meets, meaning it has freedom of movement on the lattice at no cost to acceptance probability. One can view this advantage as “searching” over a superposition of possible states of similar energy, before collapsing to one at measurement. This becomes incredibly important when in the low temperature regime. As mentioned in Section II.2, Markov-chain suffer from low acceptance rate in the low temperature regime so proposals that can suggest similar energy states are highly likely to be accepted and can significantly improve a Markov-chains convergence.
III Framework and figures of merit
In this section, we establish the foundational framework for employing CGQeMCMC techniques on the Ising model by outlining the key components and methodologies. Additionally, we provide a comprehensive overview of two pivotal figures of merit: the absolute spectral gap and magnetization estimation. These metrics are introduced to provide background for the numerical experiments detailed in Section V.
III.1 Framework
In CG modeling, the behavior of a complex system is simulated by using a simpler representation of the system. This is immediately realized in Ising models at each MCMC step by isolating a group consisting of $q$ spins - here referred to by the corresponding fraction of the lattice per group $(q/n)$. These spins may not necessarily be close to one another or even interact, however the optimal (in terms of thermalization) choice of spins in a group is likely to depend on the details of the problem. The coarse graining is shown in Algorithm 3, where the QeMCMC (Algorithm 2) is run as a subroutine on a small group $s_{g}$.
@@ 8 $s_{g}^{\prime}\leftarrow$ QeMCMC ($s_{g}$);
// update $\mathbf{s}^{\prime}$ with results from $s^{\prime}_{g}$ QeMCMC
9 $\mathbf{s}^{\prime}[index]\leftarrow s_{g}^{\prime}$ ;
@@
Of course, there is scope for a multitude of different groupings, sub-groupings and clusterings as is sometimes seen in classical literature Zhu et al. (2015); Houdayer (2001); Wolff (1989); Swendsen and Wang (1987); Wu et al. (2021); Murray (2007). These are not exhaustively discussed here, where we simply give the framework of coarse graining the Ising model lattice along with some illustrative examples that demonstrate the potential of the approach. Discussion of the choice of spins to include in the active group, as well as the parameters chosen as input to the QeMCMC, including the choice of the “reduced” Hamiltonian, are left to Section IV. It should be made clear that any Quantum enhancement found can most likely be combined with traditional MCMC methods in hybrid schemes that are likely to outpace “plain” MCMC, regardless of proposal.
III.2 Spectral Gap
As mentioned in Section II.3, estimating the thermalization time $(\tau_{\varepsilon})$ of an algorithm is not simple, as one cannot necessarily tell when the Markov chain is in the stationary state. There is also a dependence on the starting state of the algorithm meaning between successive MCMC runs thermalization can differ in steps by orders of magnitude. The thermalization time is however bounded analytically:
@@ III.2 Spectral Gap
where $\varepsilon$ is the error in total variational distance used to quantify the bounds of $(\tau_{\varepsilon})$ Levin and Peres (2017). The Spectral gap $(\delta)$ describes the slowest facet of a chains convergence and is found through the transition probability matrix, $P$. Given the eigenvalues $\{\lambda\}$ of $P$, $\delta$ is calculated:
@@
The Spectral gap is inversely proportional to the upper bound of the thermalization time (Equation 10) and thus provides an unambiguous quantification of a Markov chains convergence. Numerical calculation of $\delta$ is very inefficient however, requiring numerical integration (often by random sampling, see Appendix A) to create the $(2^{n}\times 2^{n})$ $P$ matrix before Eigen-decomposition. Due to this, calculation of the spectral gap is exponentially more demanding than an actual MCMC algorithm.
@@
The spectral gap also ignores the use of a particular chain in more sophisticated schemes mentioned in Section II.3 that rely on running multiple chains concurrently. Thus, one must be reluctant to overstate the results of spectral gap analysis.
III.3 Convergence analysis
@@
To analyze the convergence of CGQeMCMC in test cases that is currently technologically inaccessible to QeMCMC, we run Markov Chains and estimate the average magnetization and energy of the Boltzmann distribution. For low temperatures, this essentially becomes an optimization problem as only a few states contribute.
@@
As mentioned in Section II, calculating the average magnetization $\langle m\rangle_{\mu}$ is dependent on first calculating the partition function. This makes direct calculation infeasible for even relatively simple systems. The alternative is to perform the following estimation where $m(s)$ is calculated for each sampled state by Equation 3:
$\langle m\rangle_{\mu}\approx\frac{1}{N}\sum_{\mathbf{s}\in\mathcal{S}}m(\mathbf{s})$ (12)
@@
Assuming thermalization of the Markov chain, this is accurate with high probability even when $N$ is very small in comparison to the entire state space Ambegaokar and Troyer (2010); Levin and Peres (2017). One can use estimation of this observable as a figure of merit by asking the question: “How well can our Markov chain estimate the true magnetization?”. Of course this requires knowledge of the stationary distribution. Here the stationary distribution of exemplar 16, 25 and 36 spin instances are found by brute force, meaning that the update proposals can be evaluated against the true magnetization as benchmark.
@@
Specifically, this paper will look at cumulative magnetization as sampling from Markov Chains generally requires thinning (to take non-correlated samples) and burning (to ensure thermalization) which incurs hyper-parameters complicating the analysis.
@@
Crucially, in certain cases, magnetization estimation can be a flawed figure of merit. As briefly discussed in Section II.3, a chain may appear to converge to the stationary distribution, while it is actually stuck in a local minimum. If the magnetization contribution of this local minimum is similar to that of the actual magnetization, then this becomes a flawed figure of merit. Thus, alongside the magnetization calculation we analyse the convergence of average energy between multiple chains which can be compared with the exact average energy for the stationary distribution. If local minima impede convergence, this measure will display an obvious bias.
IV Coarse graining strategies
@@
In this section, we give three different coarse graining strategies: (a) a “naive” approach where one chooses a random subset of spins and ignores the rest, (b) an improved local approach where one chooses a random subset of spins while treating the remaining spins as an external field and (c) an approach that considers multiple subsets, treating each of them separately as in the earlier approaches. Each of these coarse graining strategies is - to our knowledge - novel and of increasing complexity following our scientific exploration.
@@
Before proceeding to the proposed strategies, let us highlight two properties of the QeMCMC method that were key to the advantage demonstrated in Layden et al. (2023). The first is that the Hamiltonian used (time evolution of which obtains the proposal state) includes the problem’s Hamiltonian, and with suitable choice of parameters the proposal step does not differ significantly in energy – thus ensuring proposed steps that are accepted even when being close to thermalization. The second is that, unlike classical local strategies, QeMCMC can lead to proposal step that differs (significantly) in Hamming distance from the previous step, “spanning/searching” the full space and escaping from local minima.
IV.1 Naive CG - local groups
The naive choice of coarse graining is to pick the indices of spins that are neighbors forming the set, $g_{l}=\{i_{r},i_{r+1}...i_{r+q}\}$ where $i_{r}$ is the index of a randomly selected spin $s_{r}$ that begins our “local group”.
The “locality” of the spins refer to their indices, which in some cases (for example the 1-d Ising model) also carry physical information, while in other cases (random Ising model) do not. We still use the term “local” to distinguish this coarse graining strategy from others, but this should not be misunderstood as “physically” local. Future optimal coarse graining strategies will likely not use local groupings, but we limit our analysis in this paper to local groupings for simplicity.
Once isolated, the group is updated according to the QeMCMC routine outlined in Section II.4 with $\tilde{J}$ and $\tilde{h}$ the parameters defining the “partial Hamiltonian” of the newly formed group. This new Hamiltonian represents only the selected spins, while essentially ignoring the spins not selected: